OpenStudy (anonymous):

Need to find a shaded area between y = x^2-2x and the x-axis. Thanks!

Parth (parthkohli):

Hello. The integral of a function is the area between that and the x-axis.

OpenStudy (anonymous):

I need to set the two equations to get the integers. One I know the x-values I can sub them into my def integral. BUT...what is the y=? if the shaded area ends at the x-axis?

Parth (parthkohli):

Err... I am not able to understand you.\[y = x^2 - 2x \]essentially is the same as\[f(x) = x^2 - 2x\]Also, this is not a definite integral because there are no upper or lower limits on the \(x\). So you'll just have to find an expression for the area.

Parth (parthkohli):

But I am sorry if I misunderstood your question...

Parth (parthkohli):

Can you integrate \(x^2 - 2x\)?

OpenStudy (anonymous):

2x-2?

OpenStudy (anonymous):

I think I meant to ask whether I should try to find an expression for the x-axis. But now I understand that I need only the expression that I got, and integrate it within the limits that I have...in this case, 2 and 0

OpenStudy (kc_kennylau):

This is a definite integral, and the limits are the x-intercepts. Do you remember how to find x-intercepts? :)

OpenStudy (anonymous):

I think I meant to ask whether I should try to find an expression for the x-axis. But now I understand that I need only the expression that I got, and integrate it within the limits that I have...in this case, 2 and 0\[\int\limits_{0}^{2}\]

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

then I sub the x's and subtract F(b) -F(a), right?

OpenStudy (kc_kennylau):

so that makes \(\Large\int_0^2\normalsize (x^2-2x) dx\)

OpenStudy (kc_kennylau):

yep, except \(\int(x^2-2x)dx\ne2x-2\) (That's integration not differentiation!!!)

OpenStudy (anonymous):

hmm....then I got lost...again.

OpenStudy (anonymous):

I need to differentiate?

OpenStudy (kc_kennylau):

no you need to integrate but you differentiated

OpenStudy (kc_kennylau):

\[\large\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (kc_kennylau):

Don't give up :)

OpenStudy (anonymous):

\[\frac{ x^3 }{ 3 } -x^{2}\] ??

OpenStudy (kc_kennylau):

yep sorry for late reply

OpenStudy (anonymous):

It's quite allright. Thanks for fan-ning. And for your help. =)

OpenStudy (kc_kennylau):

but you still have to plug in x :)

OpenStudy (anonymous):

Yeah, but these are the details. I now repeated the solving algorithm of these kinds of questions.

OpenStudy (kc_kennylau):

good

OpenStudy (anonymous):

answer: -1.3(3)

OpenStudy (kc_kennylau):

yep